# Explaining higher-order correlations between elliptic and triangular flow

### v1

@@ V Relations between cumulants of different orders
Figure 4: Ratios in Eqs. (15) and (17). Symbols are ALICE data, where the mixed cumulants are taken from Ref. ALICE:2021adw and $v_{2}\{4\}$ from Ref. ALICE:2016ccg , as a function of the collision centrality in Pb+Pb collisions at 5.02 TeV per nucleon pair. Horizontal lines are our theory predictions.
Eqs. (12), (13) and (14) show that to leading order, the experimental cumulants only depend on a small number of intrinsic cumulants. First of all, they involve the mean elliptic flow in the reaction plane, which can be obtained from data using the approximation $\bar{V}_{2}\approx v_{2}\{4\}$ (App. A). Since $v_{2}\{4\}$ is only measured above 5% centrality ALICE:2016ccg , we exclude the most central bin 0-5% from now on.
@@
This prediction is tested against ALICE data in Fig. 4. The experimental ratio in the left-hand side is in rough agreement with the predicted value in the right-hand side, but somewhat smaller in absolute magnitude.
@@
Eqs. (15) and (16) are rigorous mathematical results to leading order in $\cal V$}. They generalize the well-known identities $v_{2}\{4\}=v_{2}\{6\}=v_{2}\{8\}$ (Appendix A) to mixed cumulants. We therefore expect that their accuracy is comparable, at the percent level in Pb+Pb collisions at LHC energies CMS:2017glf ; ALICE:2018rtz ; CMS:2023bvg . The discrepancy between our prediction and data is an order of magnitude larger than expected in Fig. 4. We predict that agreement will be much improved if the analysis is repeated with finer centrality bins.
@@
This prediction is also in fair agreement with data, as shown in Fig. 4. The experimental ratio is again systematically smaller than our prediction in absolute magnitude.
@@
Comparison with ALICE data is displayed in Fig. 5. The agreement with our prediction is much worse than in Fig. 4. As in Fig. 4, the ratio is smaller than our prediction in absolute magnitude. It decreases monotonically as a function of centrality, unlike the ratios in Fig. 4. Our prediction is quantitatively wrong, but its order of magnitude is correct, which is a non-trivial achievement given the simplicity of our approach.
@@
Agreement with data, displayed in Fig. 5, is even worse than for Eq. (18). Unlike what was observed with the previous ratios, the left-hand side is not smaller than the right-hand side for all centralities. This is not surprising, as there is no regime where the term involving $c_{0133}$ is negligible with respect to the first term.
@@ VI Conclusions
This simplicity arises from the general properties of local density fluctuations, which are of quantum origin. The classical fluctuations of impact parameter spoil this simplicity Roubertie:2025qps ; Alqahtani:2025wan . It is therefore essential to work with fine centrality bins. Despite the wide centrality bins used by ALICE, some of our relations are in reasonable agreement with Pb+Pb data. We expect that agreement would be significantly improved with a finer binning. If there are residual nonflow correlations, accuracy will be also improved with an increased pseudorapidity coverage.
It will be important to check whether full hydrodynamic calculations with fluctuating initial conditions, carried out at fixed impact parameter, confirm the validity of our results, which are derived under the simplifying assumption of linear hydrodynamic response. Significant deviations from linear response are observed for elliptic flow above 30% centrality Niemi:2015qia ; Noronha-Hostler:2015dbi ; Hippert:2020kde ; Giacalone:2017uqx , and their effect on mixed harmonic cumulants must be quantitatively assessed. The ALICE paper ALICE:2021adw contains full hydrodynamic calculations, done with the iEBE-VISHNU model Zhao:2017yhj . These calculations are in fair agreement with experimental results, but have large error bars. This can easily be improved. Hydrodynamic calculations of $MHC(v_{2}^{2m},v_{3}^{2q})$ with $q=1$ should require modest statistics ($q=2$ and $q=3$ are more demanding because the corresponding normalized cumulants are smaller), provided that one evaluates anisotropic flow directly on the freeze-out surface Cooper:1974mv . By contrast, the iEBE-VISHNU calculation couples the hydrodynamics to a transport calculation Bass:1998ca ; Petersen:2008dd . This is the state-of-the-art approach when it comes to comparing with experimental data, but it results in larger statistical errors, induced by the Monte Carlo sampling of hadrons. It is likely that effects of the transport phase largely cancel when taking ratios as in Sec. V. It will be also essential to carry out these hydrodynamic calculations at fixed impact parameter, which is rarely done. Impact parameter fluctuations can be studied independently Samanta:2023amp ; Alqahtani:2024ejg .
If precise hydrodynamic calculations and future experimental analyses confirm our predicions, this will be a spectacular illustration of the unreasonable effectiveness of hydrodynamics Noronha-Hostler:2015wft in describing multiparticle correlations observed in nucleus-nucleus collisions.
@@ Acknowledgements.
M. Alqahtani acknowledges the support of the Research Mobility Program of the French Embassy in Riyadh, which helped to initiate this work.
@@ Appendix A Moments and cumulants of $v_{2}$
We use these equations to evaluate the denominators in Eq. (5). The cumulants in the right-hand side of Eq. (22) are measured in the same experiment with the same centrality bins ALICE:2016ccg . We neglect $v_{3}\{4\}$ and $v_{3}\{6\}$, which are not provided in this reference, and whose contribution is known to be small ATLAS:2019peb . For the 0-5% centrality bin, $v_{2}\{4\}$ and $v_{2}\{6\}$ are undefined (the corresponding cumulants change sign STAR:2015mki ; ATLAS:2019peb ; Alqahtani:2024ejg ) and we set them to zero.

### v2

@@ V Relations between cumulants of different orders
Figure 5: Ratios in Eqs. (15) and (17) as a function of the collision centrality in Pb+Pb collisions at 5.02 TeV per nucleon pair. Open symbols are ALICE data, where the mixed cumulants are taken from Ref. ALICE:2021adw and $v_{2}\{4\}$ from Ref. ALICE:2016ccg . Closed symbols are CMS data CMS:2025opi . Horizontal lines are our theory predictions, Eqs. (15) and (17).
Eqs. (12), (13) and (14) show that to leading order, the experimental cumulants only depend on a small number of intrinsic cumulants. First of all, they involve the mean elliptic flow in the reaction plane, which can be obtained from data using the approximation $\bar{V}_{2}\approx v_{2}\{4\}$ (App. A). ALICE only measures $v_{2}\{4\}$ above 5% centrality ALICE:2016ccg so that we exclude the most central bin 0-5%.
@@
Fig. 5 shows that CMS data agree very well with this equation, except for the most central bin, where impact parameter fluctuations may have a sizable effect. Agreement with ALICE data is poorer.
@@
Equations (15) and (16) are rigorous mathematical results to leading order in $\cal V$}. They generalize the well-known identities $v_{2}\{4\}=v_{2}\{6\}=v_{2}\{8\}$ (Appendix A) to mixed cumulants. We therefore expect that their accuracy is comparable, at the percent level in Pb+Pb collisions at LHC energies CMS:2017glf ; ALICE:2018rtz ; CMS:2023bvg , provided that the analysis is carried out in fine centrality bins. [^11] The intrinsic centrality resolution at the LHC is of order 2% Das:2017ned . This is the typical order of magnitude of the optimal width of the centrality binning.
@@
This prediction is also in good agreement with data, as shown in Fig. 5. Eqs. (15), (16), and (17) show that as one increases $m$, the change in $MHC(v_{2}^{2m},v_{3}^{2})$ is solely determined by the mean elliptic flow in the reaction plane $\bar{V}_{2}\simeq v_{2}\{4\}$. This property also holds for higher orders in $v_{3}$, which we now study.
@@
Comparison with experimental data is displayed in Fig. 6. Our prediction agrees well with CMS data, not with ALICE data.
@@
This prediction roughly matches CMS data in terms of order of magnitude (Fig. 6). Here, we do not expect any significant improvement with a finer centrality binning, as we have neglected the term involving $c_{0133}$ which is generically of the same order of magnitude as the other terms.
@@ VI Conclusions
This simplicity arises from the general properties of local density fluctuations, which are of quantum origin. The classical fluctuations of impact parameter spoil this simplicity Roubertie:2025qps ; Alqahtani:2025wan . It is therefore essential to work with fine centrality bins. The wide centrality bins implemented by ALICE partly explain the differences between data and our predictions. The recent CMS data are in much better agreement with our calculation. This illustrates that the effect of impact parameter fluctuations is modest. We expect that agreement would be further improved with a finer centrality binning. In addition, we have made a number of quantitative predictions for cumulants of order 10 (Eqs.(16) and (19)), which could be analyzed with existing data.
It will be important to check whether full hydrodynamic calculations with fluctuating initial conditions, carried out at fixed impact parameter, confirm the validity of our results, which are derived under the simplifying assumption of linear hydrodynamic response. Significant deviations from linear response are observed for elliptic flow above 30% centrality Niemi:2015qia ; Noronha-Hostler:2015dbi ; Hippert:2020kde ; Giacalone:2017uqx , and their effect on mixed harmonic cumulants must be quantitatively assessed. The ALICE ALICE:2021adw and CMS CMS:2025opi include comparisons of data with full hydrodynamic calculations, done with the iEBE-VISHNU model Zhao:2017yhj and with the IP-GLASMA+MUSIC+URQMD framework McDonald:2017ayb ; Gale:2013da ; Bass:1998ca . These calculations are in fair agreement with experimental results, but have large error bars. This can easily be improved. Hydrodynamic calculations of $MHC(v_{2}^{2m},v_{3}^{2q})$ with $q=1$ should require modest statistics ($q=2$ and $q=3$ are more demanding because the corresponding normalized cumulants are smaller), provided that one evaluates anisotropic flow directly on the freeze-out surface Cooper:1974mv . By contrast, both iEBE-VISHNU and IP-GLASMA+MUSIC +URQMD couple the hydrodynamics to a transport calculation Bass:1998ca ; Petersen:2008dd . This is the state-of-the-art approach when it comes to comparing with experimental data, but it results in larger statistical errors, induced by the Monte Carlo sampling of hadrons. It is likely that effects of the transport phase largely cancel when taking ratios as in Sec. V. It will be also essential to carry out these hydrodynamic calculations at fixed impact parameter, which is rarely done. Impact parameter fluctuations can be studied independently Samanta:2023amp ; Alqahtani:2024ejg , and we have argued that their effects should be small above 5-10% centrality.
@@ Acknowledgements.
We thank Ante Bilandzic and You Zhou for discussions about the ALICE results. M. Alqahtani acknowledges the support of the Research Mobility Program of the French Embassy in Riyadh, which helped to initiate this work.
@@ Appendix A Moments and cumulants of $v_{2}$
The cumulants in the right-hand side of Eq. (22) are taken from each experiment and used to evaluate the denominators in Eq. (5). For ALICE results, they are measured with the same centrality bins ALICE:2016ccg . We neglect $v_{3}\{4\}$ and $v_{3}\{6\}$, which are not provided in this reference, and whose contribution is known to be small ATLAS:2019peb . For the 0-5% centrality bin, $v_{2}\{4\}$ and $v_{2}\{6\}$ are undefined (the corresponding cumulants change sign STAR:2015mki ; ATLAS:2019peb ; Alqahtani:2024ejg ) and we set them to zero. On the other hand, for CMS results they are taken from CMS:2025opi with finer centrality bins. We note here that $v_{3}\{4\}$ is provided in this CMS analysis.